OpenStudy (moongazer):

Marian bought a pie. She gave three parts of eight to Jake. Then, 1/3of the remaining pie was given to Sheila. Lastly, she shared the half of what was left to Henry. How much is left for Marian? Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

1- 3/8 - 1/3 * (1- 3/8) - 1/2 * (1 - 3/8 - (1/3 * (1 - 3/8))) 1 - 3/8 - 1/3 * 5/8 - 1/2 * (1 - 3/8 - (1/3 * 5/8)) 1 - 3/8 - 5/24 - 1/2 * (1- 3/8 - 5/24) 1 - 3/8 - 5/24 - 1/2 * (1 - 9/24 -5/24) 1 - 3/8 - 5/24 - 1/2 * (10/24) 1 - 3/8 - 5/24 - 5/24 24/24- 9/24 - 5/24 -5/24 24/24-(9+5+5)/24 (24-19)/24 5/24

OpenStudy (moongazer):

WOW! i think that is very long solution. could you give me a simple one?

OpenStudy (anonymous):

It's written out long so you can see the steps It's also kind of in columns so you can see what gets given to each person 3/8 is given to Jake 5/24 is given to Sheila 5/24 to Harry Which leaves 5/24

OpenStudy (anonymous):

First, ignoring everything else, after she gives some to Jake and Sheila, she splits whatever remains between herself and Henry ("she shared half of what was left to Henry"). So essentially, all we need to do is determine how much she gave to Henry and she will have the same amount. So, let's look at how much everyone got. She gave 3 parts of 8 to Jake. This is the same as saying she gave 3/8 to Jake. That leaves 5/8 of the pie remaining. From this, she gave 1/3 of the remaining 5/8 to Sheila. 1/3 of 5/8 is the same as 1/3 * 5/8 = 5/24. So between Jake and Sheila, she gave away 3/8 + 5/24. Getting a common denominator of 24, we have 3/8 + 5/24 = 9/24 + 5/24 = 14/24. So 14 out of 24 pieces has been given away. That leaves 10 of 24 to be shared between Marian and Henry. Splitting the 10 pieces evenly gives 5/24 to Henry...and consequently Marian as well. Answer: 5/24

OpenStudy (moongazer):

in simple explanation why henry will have the same amount of pie i slightly understand it but not that well

OpenStudy (anonymous):

Actually, even easier to see is that after she gives the 3 pieces of 8 to Jake, she has a certain amount left. In actuallity, it's 5/8, but really the amount is not relevant. This is because she gives 1/3 of the remaining to Sheila. SHe then has 2/3 of the remaining to split between her and Henry. Half of 2/3 is 1/3 which is the same as what she gave Sheila. So Marian, Henry, and Sheila will always have the same amount regardless of how much she gave to Jake.

OpenStudy (anonymous):

Because no matter how much she gave away to Jake and Sheila, she has a particular amount remaining. The problem them states that she shared half of whatever was left to Henry. So no matter what she's left with after Jake and Sheila get theirs, She and Henry split the rest.